OpenStudy (anonymous):

2. Solve 10^6x = 93. Round to the nearest ten-thousandth. (1 point) 11.8109 1.0986 13.3801 0.3281 3. Use a graphing calculator. Solve 5.5^3x = 805 by graphing. Round to the nearest hundredth. (1 point) 2.91 3.92 1.31 0.97

OpenStudy (anonymous):

1 (Degrees/Real Numbers) Input solve⁡(〖10〗^6 x=93,x) Solution 1 x=93/1000000≈9.3•〖10〗^(-5) 2 (Degrees/Real Numbers) Input solve⁡((5.5)^3 x=805,x) Solution 1 x=6440/1331≈4.8384673178062 You should check out microsoft mathematices!

OpenStudy (anonymous):

Sorry, have that first one off. 3 (Degrees/Real Numbers) Input solve⁡(〖10〗^(6 x)=93,x) Solution 1 x=log⁡(93)/6≈0.3280804914257

OpenStudy (anonymous):

For the record, the answer to the second one was 1.31